### v1

# A Novel Hierarchical Co-Optimization Framework for Coordinated Task Scheduling and Power Dispatch in Computing Power Networks

Abstract
The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which promise to deliver ubiquitous and on-demand computational resources. However, the immense energy consumption of these networks poses a significant sustainability challenge. Simultaneously, power grids are grappling with the instability introduced by the high penetration of intermittent renewable energy sources (RES). This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations. The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment (SUC) stage and a real-time operational stage. The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a Deep Reinforcement Learning (DRL) agent. This agent makes intelligent, carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity. Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches. Results demonstrate that the proposed framework reduces total carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining stringent Quality of Service (QoS) for computational tasks.
Index Terms:
Computing Power Network (CPN), renewable energy, data center, carbon-aware scheduling, deep reinforcement learning.
I Introduction
I-A Background
The digital transformation of the global economy is fueling an unprecedented demand for computational power. The rise of artificial intelligence (AI), big data analytics, and the Internet of Things (IoT) has pushed traditional centralized cloud computing architectures to their limits [1]. In response, a new paradigm known as the Computing Power Network (CPN) has emerged [2]. A CPN aims to interconnect vast, geographically distributed, and heterogeneous computing resources, spanning from large-scale data centers to edge nodes, into a unified, programmable fabric [3]. The core vision of CPN is to break down resource silos and enable the flexible, on-demand scheduling and allocation of computing, storage, and network resources, thereby improving efficiency and enabling novel, low-latency applications.
Concurrently, the global energy sector is undergoing a profound transition towards sustainability, marked by massive investments in Renewable Energy Sources (RES) such as wind and solar power [4]. This shift presents a powerful, yet challenging, symbiotic relationship with the burgeoning CPN paradigm. On one hand, CPNs are voracious consumers of electricity; data centers alone are projected to account for a significant portion of global electricity load growth, with AI workloads being a primary driver. Globally, data centers already account for $1-2\%$ of total electricity consumption, a figure comparable to the aviation industry. Driven by the explosive growth of AI, this demand is projected to double by 2030 [5]. In the United States, data center electricity usage is forecast to climb from $4.4\%$ of the national total in 2023 to as high as $12\%$ by 2028111https://www.energy.gov/articles/doe-releases-new-report-evaluating-increase-electricity-demand-data-centers, with AI workloads shifting from a minor component to a primary driver of this expansion. The availability of clean, renewable energy offers a direct path to decarbonize this massive computational infrastructure[6]. On the other hand, the very nature of CPNs, with their inherent flexibility in workload scheduling, presents a unique opportunity to support and stabilize a grid increasingly reliant on intermittent RES [7].
I-B Research Motivation
The central challenge lies in a fundamental misalignment between the operational dynamics of CPNs and RES-dominated power grids. CPNs are designed to host a wide array of computational tasks, from latency-sensitive services like cloud virtual reality to compute-intensive, batch-processing jobs like large model training [8], [9]. These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS). However, RES are inherently variable, non-dispatchable, and dependent on weather conditions. This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10], [11].
Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict. During periods of low RES output, the CPN’s demand would force the grid to rely on expensive and carbon-intensive fossil-fuel peaker plants to maintain balance. Conversely, during periods of high RES output, the grid may be forced to curtail wind or solar generation to prevent overload, even as CPNs continue to draw power from a mixed-carbon source. This decoupled operation leads to a suboptimal outcome: either unreliable computation or high operational costs and carbon emissions. The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset. The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [12]. This reframes the problem from ”how to power the CPN cleanly” to ”how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”. This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
I-C Contributions of the Paper
This paper proposes a holistic, hierarchical co-optimization framework that treats the power system and the CPN as a single, integrated entity. By breaking down the silos between power system operators and CPN schedulers, the framework unlocks significant economic and environmental benefits. It coordinates power generation dispatch with computational task scheduling across multiple timescales to align CPN energy demand with the availability of low-cost, low-carbon renewable energy. As far as we know, this is one of the first efforts to jointly optimize the power grid and the CPN. The main contributions of this work are summarized as follows:
  • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid. The model incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
  • A Novel Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages. A day-ahead planning stage addresses long-term unit commitment and resource reservation, while a real-time operational stage handles dynamic economic dispatch and adaptive task scheduling.
  • Scalable Optimization with Benders Decomposition: To address the large-scale, mixed-integer nature of the day-ahead Stochastic Unit Commitment (SUC) problem, Benders decomposition is employed. This technique effectively decouples the integer commitment decisions from the continuous dispatch variables, ensuring the problem remains computationally tractable even for large systems and numerous uncertainty scenarios.
  • Adaptive Real-Time Scheduling with Deep Reinforcement Learning (DRL): A DRL agent is developed for the real-time CPN task scheduling subproblem. This model-free approach enables fast, adaptive, and carbon-aware scheduling decisions in response to the highly dynamic and complex state of the joint CPN-grid system, complementing the model-based optimization of the power dispatch.
  • Extensive High-Fidelity Performance Evaluation: The efficacy of the proposed TSCO framework is validated through extensive simulations using realistic data. The evaluation demonstrates significant improvements in carbon emissions, operational costs, and RES utilization when compared against decoupled and carbon-agnostic baseline strategies.
I-D Paper Structure
The remainder of this paper is structured as follows. Section II provides a critical review of related work in CPN scheduling, integrated energy systems, and carbon-aware computing. Section III presents the detailed mathematical formulation of the CPN and power system models, defining the joint optimization problem. Section IV describes the proposed TSCO framework, including the Benders decomposition algorithm and the DRL-based scheduler. Section V details the simulation setup and presents a comprehensive performance evaluation against several baseline methods. Finally, Section VI concludes the paper with a summary of findings and directions for future research.
II Related Works
This section provides a structured review of the literature across four key domains that intersect in this work, Computing Power Network (CPN) architectures, integrated energy system management, carbon-aware workload scheduling, and optimization under uncertainty in power systems. This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
II-A Computing Power Network Architectures and Scheduling
The concept of the CPN has evolved from earlier paradigms like cloud, fog, and edge computing, with the primary goal of creating a unified network for ubiquitous computing resources. Early research focused on defining the architecture and core features of CPNs, such as intent-driven operation, closed-loop autonomy, and elastic scheduling. Architectures have been proposed with both centralized control planes, which possess a global view for unified scheduling, and distributed schemes, where decisions are made locally by network nodes.
A significant body of research in CPNs has concentrated on task scheduling [13]. The primary objectives have traditionally been to optimize QoS metrics. For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques. Other works have explored task offloading in terminal-side CPNs or the joint selection of routing paths and computing nodes [14]. While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored. Existing CPN scheduling literature typically treats the power grid as an external, static entity, overlooking the potential for dynamic, symbiotic interaction.
II-B Integrated Energy System Management
In the power systems domain, the concept of coordinating generation, transmission, and consumption has been studied extensively under the ”Generation-Grid-Load-Storage” integrated operation model [11]. This paradigm seeks to improve system safety, economy, and reliability through the coordinated interaction of all system components. Research in this area includes the development of multi-timescale optimal dispatching strategies [15], economic dispatch models that incorporate DR, and optimal power flow (OPF) formulations that aim to minimize generation costs while respecting network constraints [16].
The integration of flexible loads and DR has been identified as a key enabler for grids with high RES penetration. Studies have explored how to coordinate data centers as flexible loads with a load aggregator to minimize electricity costs and absorb grid volatility [17]. However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload. They treat the CPN as a ”black box” load, missing the opportunity to optimize its internal operations in concert with the grid.
II-C Carbon-Aware Workload Scheduling
The field of green computing has produced a significant body of work on carbon-aware scheduling for geo-distributed data centers. A common strategy is spatio-temporal scheduling [18], which involves shifting computational workloads in time (”temporal flexibility”) or space (”geographical flexibility”) to data centers with lower electricity prices or cleaner energy mixes. These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [19]. This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a ”virtual queue” that tracks the deviation from the budget. The scheduler is then penalized for actions that increase this queue length [20]. While powerful, this line of research suffers from a critical limitation. It almost universally treats the power grid as an exogenous system. The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions. This assumption breaks down at scale, as the collective actions of large CPNs can and will influence grid operations, market prices, and the generation mix, thereby altering the very signals they are responding to.
II-D Research Gap Summary
The existing body of work, while extensive in its respective domains, reveals a significant research gap at the intersection of CPNs and power systems. Current research either:
  • Simplifies the power grid: CPN and carbon-aware scheduling studies treat the grid as a static source of price and carbon signals, ignoring the feedback loop where scheduling decisions impact the grid.
  • Simplifies the CPN: Power system and DR studies model flexible loads like data centers in an overly simplistic manner, failing to capture the rich internal complexity of CPN workloads, resource heterogeneity, and QoS constraints.
This leads to a new class of control problem where the decision variables are distributed across two separate domains. The system dynamics are characterized by a mix of well-understood physics (the power grid) and complex, stochastic behavior (the CPN). A model-based optimization approach is ill-suited for the fast, dynamic CPN scheduling, while a model-free AI approach cannot guarantee adherence to the hard physical constraints of the power grid. Consequently, there is a clear need for a hybrid framework that endogenously models the bi-directional interactions between the two systems. Our TSCO framework, which combines large-scale, model-based optimization for slow, physics-heavy planning with a model-free, adaptive AI technique for fast, complex real-time scheduling, is designed specifically to fill this gap.
III System Architecture and Problem Formulation
This section presents the mathematical models for the CPN and the integrated power system. These models form the basis of our co-optimization framework, as shown in Fig. 1.
Figure 1: CPN and power grid co-optimization architecture.
III-A Computing Power Network (CPN) Model
The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, representing data centers, and $\mathcal{L}$ is the set of communication links connecting them.
III-A1 CPN Node Model
Each CPN node $n\in\mathcal{N}$ is characterized by its available computational resources, such as CPUs, GPUs, and TPUs. For simplicity in this formulation, we aggregate these into a single measure of computing capacity, $C_{n}^{comp}$, expressed in Floating Point Operations Per Second (FLOPS). The power consumption of a CPN node is modeled as a linear function of its utilization [21], varying between an idle power $P_{n}^{idle}$ and a peak power $P_{n}^{peak}$ at full utilization. The total power consumed by node $n$ at time $t$ is a decision variable, $P_{n,t}^{CPN}$, determined by the scheduled tasks.
III-A2 Task Model
We model incoming computational jobs as Directed Acyclic Graphs (DAGs), a common representation for parallel applications with precedence constraints [22], [23]. A job $k\in\mathcal{J}$ is represented by $J_{k}=(\mathcal{T}_{k},\mathcal{E}_{k})$, where $\mathcal{T}_{k}$ is the set of sub-tasks and $\mathcal{E}_{k}$ is the set of directed edges representing dependencies. An edge $(\tau_{i},\tau_{j})\in\mathcal{E}_{k}$ implies that sub-task $\tau_{j}$ cannot begin until $\tau_{i}$ is complete. Each sub-task $\tau\in\mathcal{T}_{k}$ is defined by its total computational workload $w_{\tau}$ (in floating-point operations) and its resource requirement $r_{\tau}$ (e.g., number of processing units). The execution time of sub-task $\tau$ on node $n$ is thus $t_{\tau,n}=w_{\tau}/C_{n}^{comp}$. Each job $k$ has an arrival time $A_{k}$ and a hard end-to-end deadline $D_{k}$.
III-B Integrated Power System Model
The power grid is modeled based on the IEEE 30-bus test system [^3] https://icseg.iti.illinois.edu/ieee-30-bus-system/, a standard benchmark for power system studies. The system consists of a set of buses $\mathcal{I}$ connected by transmission lines. Each CPN node $n\in\mathcal{N}$ is co-located with a specific load bus $i\in\mathcal{I}$.
III-B1 Conventional Generation
The set of conventional thermal generators, $\mathcal{G}_{C}$, forms the dispatchable backbone of the system. The fuel cost of each generator $g\in\mathcal{G}_{C}$ is represented by a quadratic function of its power output $P_{g,t}$:
$C_{g}(P_{g,t})=a_{g}P_{g,t}^{2}+b_{g}P_{g,t}+c_{g},$ (1)
where $a_{g},b_{g},c_{g}$ are cost coefficients. These generators are subject to operational constraints, including minimum and maximum power output limits ($P_{g}^{min},P_{g}^{max}$), and ramp-up/ramp-down rate limits ($RU_{g},RD_{g}$) that constrain how quickly their output can change between time periods [24].
III-B2 RES
The set of RES generators, $\mathcal{G}_{R}$, includes wind and solar farms. Their power output is non-dispatchable and uncertain. We model their available power at time $t$ in scenario $\omega$, $P_{g,t,\omega}^{R,avail}$, as a stochastic parameter derived from historical weather data. The actual dispatched power $P_{g,t,\omega}$ can be less than or equal to the available power.
III-B3 BESS
BESS units, located at specific buses, provide crucial flexibility for managing RES intermittency. Each BESS $b\in\mathcal{B}$ is modeled by its state-of-charge (SOC) dynamics [25]:
$SOC_{b,t,\omega}=SOC_{b,t-1,\omega}+(\eta_{b}^{c}P_{b,t,\omega}^{chg}-\frac{1}{\eta_{b}^{d}}P_{b,t,\omega}^{dis})\Delta t.$ (2)
The model is subject to constraints on the SOC level ($SOC_{b}^{min}\leq SOC_{b,t,\omega}\leq SOC_{b}^{max}$), where $E_{b}^{max}$ is the energy capacity, and maximum charging/discharging power ($P_{b}^{c,max},P_{b}^{d,max}$).
III-C Uncertainty and Carbon Modeling
III-C1 Uncertainty Modeling
The uncertainties in RES generation and CPN job arrivals are critical to the problem. We adopt a scenario-based stochastic programming approach [26]. A set of discrete scenarios $\Omega$ is generated, where each scenario $\omega\in\Omega$ represents a plausible joint realization of RES power availability and CPN workload over the time horizon $T$. Each scenario is assigned a probability $\pi_{\omega}$, with $\sum_{\omega\in\Omega}\pi_{\omega}=1$.
III-C2 Carbon Intensity Modeling
The environmental impact is quantified through carbon emissions. The carbon intensity of the grid is not static but depends on the real-time generation mix. The total carbon emission rate at time $t$ in scenario $\omega$, $E_{t,\omega}$, is calculated as the sum of emissions from all active generators:
$E_{t,\omega}=\sum_{g\in\mathcal{G}_{C}}\epsilon_{g}P_{g,t,\omega},$ (3)
where $\epsilon_{g}$ is the emission factor (e.g., in tons of $CO_{2}$ per MWh) of generator $g$. For RES, $\epsilon_{g}$ is zero. This endogenous calculation is crucial, as it directly links dispatch decisions to carbon output. We also leverage real-world marginal carbon intensity data, such as that provided by WattTime, to inform the real-time DRL agent about the emissions impact of consuming an additional unit of electricity at a specific location and time.
III-D Joint Optimization Problem Formulation
The overarching goal is to co-optimize the power system operation and CPN task scheduling to minimize the total expected system cost over a planning horizon $T$. The total cost comprises the operational costs of the power system and the monetized cost of carbon emissions. The problem is formulated as a large-scale, two-stage stochastic mixed-integer linear program (MILP).
Objective Function:
$\min\sum_{\omega\in\Omega}\pi_{\omega}\sum_{t=1}^{T}\left(\sum_{g\in\mathcal{G}_{C}}+\lambda_{CO_{2}}E_{t,\omega}\right).$ (4)
The objective minimizes the expected sum of three components across all scenarios: (1) the quadratic fuel costs of conventional generators, (2) the costs associated with starting up ($SU_{g}$) and shutting down ($SD_{g}$) these generators based on their commitment status $u_{g,t}$, and (3) a carbon tax, where $\lambda_{CO2}$ is the price of carbon and $E_{t,\omega}$ is the total emissions.
Key Constraints: The optimization is subject to a comprehensive set of constraints that couple the two systems:
  • Power System Constraints (for each $t,\omega$):
    Power Balance (DC-OPF): At each bus $i\in\mathcal{I}$, the total power injected must equal the total power withdrawn. This is the core DC power flow equation [27].
    $\sum_{g\in\mathcal{G}(i)}P_{g,t,\omega}+\sum_{b\in\mathcal{B}(i)}(P_{b,t,\omega}^{dis}-P_{b,t,\omega}^{chg})\\ -(P_{i,t}^{D}+P_{i,t,\omega}^{CPN})=\sum_{j\in\mathcal{I}}B_{ij}(\theta_{i,t,\omega}-\theta_{j,t,\omega}),$ (5)
    where, $\mathcal{G}(i)$ and $\mathcal{B}(i)$ are generators and BESS at bus $i$, $P_{i,t,\omega}^{CPN}$ is the CPN power demand at that bus, and the right-hand side represents the net power flow out of the bus.
    Transmission Line Limits: The power flow $F_{ij}$ on each line $(i,j)$ must not exceed its thermal limit $F_{ij}^{max}$.
    $F_{ij}^{max}\leq B_{ij}(\theta_{i,t,\omega}-\theta_{j,t,\omega})\leq F_{ij}^{max}.$ (6)
    Generator Constraints: Including commitment logic, min/max output, and ramping limits for all $g\in\mathcal{G}_{C}$.
    BESS Constraints: Including SOC dynamics, capacity limits, and charge/discharge power limits for all $b\in\mathcal{B}$.
  • CPN Task Scheduling Constraints (for each $t,\omega$):
    Task Assignment: Each sub-task $\tau$ of each job $k$ must be scheduled exactly once.
    $\sum_{n\in\mathcal{N}}\sum_{t=A_{k}}^{D_{k}}x_{k,\tau,n,t}=1\quad\forall k,\tau\in\mathcal{T}_{k}.$ (7)
    Precedence Constraints: For any dependency $(\tau_{i},\tau_{j})\in\mathcal{E}_{k}$, the start time of $\tau_{j}$ must be after the finish time of $\tau_{i}$.
    Deadline Satisfaction: The completion time of the final sub-task of job $k$ must be no later than its deadline $D_{k}$.
    Node Resource Capacity: The total resource demand of tasks scheduled on node $n$ at time $t$ cannot exceed its capacity.
    $\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau}x_{k,\tau,n,t}\leq C_{n}^{comp}\quad\forall n,t.$ (8)
  • Coupling Constraint: This is the critical link. The power consumed by a CPN node $n$ (co-located at bus $i$) is determined by the tasks scheduled on it.
    $P_{i,t,\omega}^{CPN}=P_{n}^{idle}+(P_{n}^{peak}-P_{n}^{idle})\frac{\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}}{C_{n}^{comp}}.$ (9)
    This equation makes the CPN load term $P_{i,t,\omega}^{CPN}$ in the power balance equation an endogenous variable dependent on the scheduling decisions $x_{k,\tau,n,t}$.
  • Carbon Budget Constraint: A long-term constraint on total carbon emissions is imposed to ensure sustainability goals are met.
    $\sum_{\omega\in\Omega}\pi_{\omega}\sum_{t=1}^{T}E_{t,\omega}\leq E_{budget}.$ (10)
    Due to its long-term nature, this constraint is difficult to handle directly in a short-term optimization. It will be managed implicitly through the design of the DRL agent’s reward function, as detailed in the next section.
IV Two-Stage Co-Optimization (TSCO) Framework
The joint optimization problem formulated in Section III is a large-scale, non-convex, mixed-integer stochastic program, which is computationally intractable to solve directly [28]. To address this challenge, we propose a Two-Stage Co-Optimization (TSCO) framework that decomposes the problem by decision timescale and complexity. The framework combines model-based optimization for long-term, system-wide planning with a model-free, AI-based approach for fast, adaptive real-time control, as shown in Fig. 2.
Figure 2: Two-Stage Co-Optimization (TSCO) framework for CPN and power grid collaborative optimization.
IV-A Hierarchical Stochastic Optimization Structure
The TSCO framework decomposes the problem into two distinct stages, reflecting the natural hierarchy of power system operations:
  • Stage 1 (Day-Ahead Planning): This stage solves an SUC problem for the upcoming 24-hour horizon. It makes the ”here-and-now” decisions, which are binding across all potential future scenarios. These decisions include the commitment (on/off) status $u_{g,t}$ of conventional generators and high-level energy reservation for BESS and CPN workload classes. The objective is to minimize the total expected cost over all scenarios, setting the operational envelope for the next stage. This stage is computationally intensive but is performed only once per day.
  • Stage 2 (Real-Time Operation): This stage operates at a much faster timescale (e.g., 5-15 minute intervals) and makes ”wait-and-see” recourse decisions as uncertainty unfolds. As the actual RES generation and CPN task arrivals are revealed, this stage executes two parallel, tightly coupled processes:
    1. 1.
      Real-Time Economic Dispatch (ED): Solves for the optimal power output of the committed generators and the charge/discharge schedule of BESS to meet the actual load at minimum cost, while respecting all grid constraints.
    2. 2.
      Real-Time CPN Task Scheduling: A DRL agent makes granular, second-by-second scheduling decisions, assigning individual tasks to specific resources within the CPN nodes. Its decisions are informed by the real-time grid state, namely prices and carbon intensity, provided by the ED.
This hierarchical structure allows the framework to be both economically optimal from a long-term planning perspective and highly adaptive to short-term dynamics.
IV-B Benders Decomposition for the Day-Ahead SUC
The day-ahead SUC problem is a large-scale MILP due to the combination of binary commitment variables and a large number of scenarios representing RES uncertainty. To solve it efficiently, we employ Benders decomposition [29], a classic technique for problems with this structure. The method iteratively decomposes the problem into a simpler master problem and a set of independent subproblems.
  • Master Problem: The master problem determines the integer variables in the first stage, the unit commitment schedules $\{u_{g,t}\}$ for all $g\in\mathcal{G}_{C}$ over the horizon $T$. It is a pure integer program that minimizes the sum of startup/shutdown costs and an estimated future cost, $\theta$, which represents the expected operational cost from the subproblems.
    $\min\sum_{t=1}^{T}\sum_{g\in\mathcal{G}_{C}}(SU_{g}(u_{g,t})+SD_{g}(u_{g,t}))+\theta.$ (11)
    Subject to: Generator minimum up/down time constraints; Benders optimality and feasibility cuts (added iteratively).
  • Subproblems: For a given commitment schedule $\{\bar{u}_{g,t}\}$ provided by the master problem, a separate, continuous linear program (LP) is solved for each scenario $\omega\in\Omega$. Each subproblem represents the economic dispatch problem for that scenario, minimizing the fuel and carbon costs subject to grid constraints.
    $\min\sum_{t=1}^{T}\left(\sum_{g\in\mathcal{G}_{C}}C_{g}(P_{g,t,\omega})+\lambda_{CO2}E_{t,\omega}\right).$ (12)
    Subject to: Power balance, line limits, generator output limits (for committed units), BESS constraints.
  • Algorithm Flow and Cut Generation: The algorithm proceeds iteratively:
    1. 1.
      The master problem is solved to obtain a candidate commitment schedule.
    2. 2.
      This schedule is passed to the subproblems, which are solved in parallel for all scenarios.
    3. 3.
      If any subproblem is infeasible (i.e., the commitment schedule cannot satisfy the load), its dual rays are used to construct a Benders feasibility cut, which is added to the master problem to exclude this infeasible solution.
    4. 4.
      If all subproblems are feasible, their optimal dual variable values are used to construct a Benders optimality cut. This cut is a linear inequality that provides a lower bound on the recourse cost $\theta$ and is added to the master problem.
    5. 5.
      The process repeats until the lower bound from the master problem and the upper bound from the subproblems converge within a specified tolerance.
IV-C DRL-based Real-Time CPN Task Scheduling
While the SUC/ED provides an economically optimal power dispatch plan, it is far too slow for the dynamic, fine-grained scheduling required within the CPN. For this, we propose a DRL-based approach. A DRL agent can learn a complex scheduling policy through interaction with the environment [30], enabling it to make near-instantaneous decisions that are adaptive to both CPN and grid conditions.
Markov Decision Process (MDP) Formulation: The CPN scheduling problem is formulated as an MDP defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$:
  • State ($s_{t}\in\mathcal{S}$): The state provides a comprehensive snapshot of the entire system at time $t$. It is a high-dimensional vector including:
    1. 1.
      CPN State: Characteristics of tasks in the queue (e.g., resource requirements, deadlines), current resource utilization and power consumption of each CPN node.
    2. 2.
      Grid State: Real-time RES generation levels, BESS state-of-charge, and crucially, the real-time locational marginal price (LMP) and marginal carbon intensity (MCI) for each bus hosting a CPN node. These signals are provided by the real-time ED solution.
  • Action ($a_{t}\in\mathcal{A}$): For the task at the head of the queue, the agent selects an action from a discrete set. An action is a tuple $(n,\tau_{type})$ representing the decision to assign the task to node $n$ to be processed by resource type $\tau_{type}$ (e.g., CPU, GPU). The action space also includes deferring the task.
  • Reward ($R_{t}\in\mathcal{R}$): The reward function is carefully designed to guide the agent towards the overall optimization objective. It is a weighted sum of multiple components:
    $R_{t}=$ $w_{rev}\cdot\text{Revenue}_{t}-w_{cost}\cdot\text{Cost}_{t}$ (13)
    $-w_{carb}\cdot\text{Carbon}_{t}-w_{pen}\cdot\text{Penalty}_{t},$
    where $\text{Revenue}_{t}$ is a positive reward for successfully completing a job; $\text{Cost}_{t}$ is the electricity cost of executing the scheduled task, calculated as $P_{n,t}^{CPN}\times\text{LMP}_{i,t}$; $\text{Carbon}_{t}$ denotes the carbon cost, calculated as $P_{n,t}^{CPN}\times\text{MCI}_{i,t}$. To enforce the long-term budget $E_{budget}$, this term is augmented using the Lyapunov optimization technique. A virtual carbon queue $Q_{t}$ is maintained, updating as
    $Q_{t+1}=\max(0,Q_{t}+\text{Carbon}_{t}-E_{budget}/T).$ (14)
    The reward is then penalized by an additional term proportional to $Q_{t}\times\text{Carbon}_{t}$, which strongly discourages carbon-intensive actions when the system is already over its carbon budget. And the $\text{Penalty}_{t}$ represents a large negative penalty for missing a task’s deadline.
Given the large, continuous state space and discrete action space, a value-based DRL algorithm such as Deep Q-Network (DQN) or its advanced variants (e.g., Dueling DQN, Rainbow) is suitable [31]. The agent’s policy $\pi(a_{t}|s_{t})$ is represented by a deep neural network that approximates the optimal action-value function $Q^{*}(s,a)$. The agent is trained offline on a rich dataset of historical system states and transitions, and then deployed for fast online inference.
IV-D Overall TSCO Algorithm and Its Complexity Analysis
The complete operational flow of the TSCO framework integrates the day-ahead planning and real-time control stages. The step-by-step procedure is outlined in Alg. 1.
Input : Set of RES/CPN scenarios $\Omega$ with probabilities $\pi_{\omega}$
Output : 24-hour unit commitment $\{u_{g,t}^{*}\}$, Real-time power dispatch, Real-time CPN task schedule
Stage 1: Day-Ahead SUC (solved once daily):
1. Initialize Benders master problem with generator constraints
2. repeat
   3. Solve MILP master problem to get candidate commitment $\{\bar{u}_{g,t}\}$
   4. for each scenario $\omega\in\Omega$ in parallel do
      5. Solve LP dispatch subproblem with fixed commitments $\{\bar{u}_{g,t}\}$
      6. if subproblem is infeasible then
         Generate and add a feasibility cut to the master problem
    else
         Generate and add an optimality cut to the master problem
until lower and upper bounds converge;
7. Obtain final 24-hour unit commitment schedule $\{u_{g,t}^{*}\}$
Stage 2: Real-Time Operation (for $t=1,\dots,T$):
1. for $t=1,\dots,T$ do
   2. Observe actual RES generation $P_{g,t}^{R,actual}$ and new CPN job arrivals
   3. Update CPN task queue
   4. Solve real-time Economic Dispatch for $\{u_{g,t}^{*}\}$ and BESS
   5. Obtain real-time LMPs and MCIs for all CPN node locations
   6. Construct state vector $s_{t}\leftarrow$ (CPN state + Grid state)
   7. DRL agent takes action $a_{t}\leftarrow\pi(s_{t})$ to schedule next CPN task
   8. Update CPN power demand $P_{i,t}^{CPN}$ based on action $a_{t}$
   9. Execute scheduled task and power dispatch
   10. Update BESS SOC and CPN resource status
Algorithm 1 Two-Stage Co-Optimization (TSCO) Framework
Additionally, the computational complexity of the TSCO framework is best analyzed by examining its two stages separately.
IV-D1 Stage 1: Day-Ahead SUC
The SUC problem is a mixed-integer programming (MIP) problem, which is NP-hard. Solving the full extensive form directly is computationally prohibitive for realistic system sizes and a large number of scenarios. Benders decomposition is employed to manage this complexity.
  • Master Problem: The master problem is a MILP. Its complexity is, in the worst case, exponential in the number of integer variables, which is proportional to the number of conventional generators and the length of the time horizon ($|\mathcal{G}_{C}|\times T$). The size of the master problem also grows with each iteration as Benders cuts are added.
  • Subproblems: For each of the $|\Omega|$ scenarios, a linear program (LP) is solved. The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints. Since the subproblems are independent for a given commitment schedule, they can be solved in parallel. The time taken per iteration for this step is thus equivalent to solving a single LP.
IV-D2 Stage 2: Real-Time Operation
This stage must operate quickly at each time step $t$.
  • Real-Time Economic Dispatch: This is a standard LP, similar in structure to a Benders subproblem but for a single realized scenario. As an LP, it can be solved very efficiently in polynomial time, which is essential for real-time control.
  • DRL-based CPN Scheduling: The online decision-making process involves a single forward pass through the trained deep neural network. The complexity of a forward pass is approximately $O(\sum_{l=1}^{L}N_{l}\times N_{l-1})$, where $L$ is the number of layers and $N_{l}$ is the number of neurons in layer $l$. This computation is extremely fast and independent of the complexity of the underlying system dynamics, making it highly suitable for real-time, low-latency scheduling decisions. Also, the computationally intensive training of the DRL agent is performed offline and does not impact the online operational complexity.
In summary, the TSCO framework strategically manages computational complexity by solving the NP-hard, large-scale planning problem (SUC) offline on a day-ahead basis, where longer computation times are acceptable. It then leverages highly efficient, polynomial-time algorithms (LP for ED) and fast neural network inference (DRL for scheduling) for the real-time operational stage, ensuring the framework is viable for practical deployment.
V Performance Evaluation
This section presents a comprehensive empirical validation of the proposed Two-Stage Co-Optimization (TSCO) framework. A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS.
V-A Simulation Setup
The simulation framework is implemented in Python. The power system dynamics are modeled using PyPSA, a powerful open-source library for power system analysis. The CPN and the scheduling logic are implemented as a custom discrete-event simulator. The DRL agent is developed using PyTorch. For solving the MILP and LP problems in the Benders decomposition, we use the Gurobi optimizer.
The simulation is based on a modified IEEE 30-bus test system. It includes 6 conventional thermal generators, 4 utility-scale BESS units, and 5 large-scale renewable generation sites (3 solar, 2 wind).To ensure realism, we use real-world time-series data to model the stochastic RES generation. Solar irradiance data for locations in California is sourced from the National Renewable Energy Laboratory’s (NREL) National Solar Radiation Database (NSRDB) [^4] https://catalog.data.gov/dataset/national-solar-radiation-database-nsrdb. Wind power generation profiles for locations in Germany are obtained from the ENTSO-E Transparency Platform [^5] https://www.entsoe.eu/data/power-stats/. These datasets are used to generate 100 distinct 24-hour scenarios for the SUC problem.
Figure 3: Baseline performance comparison. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
The CPN consists of 5 geo-distributed nodes, each co-located with a major load bus in the IEEE 30-bus system. The arrival patterns and resource requirements of computational jobs are derived from processed Google Cluster Data traces [^6] https://www.kaggle.com/datasets/derrickmwiti/google-2019-cluster-sample, which provide a realistic representation of large-scale data center workloads. The precedence constraints and dependency structures within jobs are modeled based on common scientific workflow patterns, such as pipeline workflow, available from the Pegasus Workflow Management System [^7] https://pegasus.isi.edu/documentation/examples/.
To ground our environmental and economic calculations in reality, we incorporate two external data sources. Real-time marginal carbon intensity (MCI) data is obtained via the WattTime API, which provides 5-minute resolution data on the emissions impact of consuming an additional MWh of electricity in various grid regions [^8] https://watttime.org/data-science/data-signals/marginal-co2/. Historical hourly locational marginal price (LMP) data is sourced from the California Independent System Operator (CAISO) public database [^9] https://www.gridstatus.io/live/caiso.
V-B Comparison Schemes
To rigorously evaluate the performance of our TSCO framework, we compare it against three baseline methods that represent alternative approaches to the problem:
  • Cost-Only Optimizer (CO-Opt) [32]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero. This represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without explicit consideration for environmental impact.
  • Renewable-Greedy Scheduler (RG-Sched) [33]: This is a heuristic-based CPN scheduling approach where tasks are always dispatched to the CPN node with the highest instantaneous RES power availability. The power system is not co-optimized; it simply reacts to the resulting CPN load profile. This baseline tests the efficacy of a simple “follow the renewables” strategy that ignores grid constraints and economic signals.
  • Decoupled Framework (DC-Frame) [34]: This baseline represents the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately. The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities. Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals. This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
V-C Simulation Results
Figure 4: Sensitivity analysis with varying carbon price. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
V-C1 Baseline Performance Comparison
In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of \$50\\$50/ton. Each data point presented is the result of $10$ independent simulation runs, from which we calculate the mean and standard deviation and present in Fig. 3.
TSCO simultaneously achieves the lowest total operational cost and the lowest carbon emissions. The CO-Opt baseline, being carbon-agnostic, minimizes only direct fuel costs, resulting in $41.5\%$ higher emissions and $17.5\%$ higher total costs once the carbon price is factored in. This is because it relies heavily on the cheapest available thermal generators, regardless of their carbon intensity. The RG-Sched heuristic, while intuitive, performs poorly across the board. By myopically chasing renewables, it ignores grid congestion and the economic cost of dispatching thermal generators to support its decisions, leading to the highest operational cost and only modest emission reductions. The DC-Frame performs better than the naive baselines but is still significantly outperformed by TSCO. Its reliance on static, day-ahead signals prevents it from adapting to real-time deviations between forecasted and actual grid conditions, leading to $12.7\%$ higher costs and $16.2\%$ higher emissions. A key finding is TSCO’s ability to significantly reduce RES curtailment. By treating the CPN as a flexible load that can absorb surplus renewable generation in real-time, TSCO reduces curtailment by over $60\%$ compared to DC-Frame and CO-Opt. This demonstrates the value of co-optimization in turning a major energy consumer into a valuable grid-stabilizing asset.
Additionally, the economic and environmental gains from TSCO do not come at the expense of computational performance. TSCO maintains a high job success rate $98.5\%$ and low average tardiness $12.3$ s, nearly on par with the cost-only optimizer. This is a direct result of the DRL agent’s reward function, which is designed to penalize deadline violations, forcing it to learn a policy that balances sustainability goals with QoS requirements. In contrast, the RG-Sched baseline suffers from a poor job success rate $85.2\%$ and high tardiness because its singular focus on RES availability often leads it to schedule tasks on nodes that are already congested, highlighting the need for a holistic system view.
V-C2 Sensitivity to Carbon Price
To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from \$25\\$25/ton to \$150\\$150/ton. This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design. The comprehensive results are presented in Fig. 4.
The TSCO framework consistently achieves a better trade-off. For any given carbon emission level, it operates at a lower cost than the DC-Frame. As the carbon price increases, both methods are incentivized to reduce emissions, but TSCO does so more efficiently. This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points. The RG-Sched method is shown as fixed points, as their operational strategy does not adapt to the carbon price.
It also shows how RES curtailment is affected by the carbon price. For both carbon-aware methods, increasing the price of carbon incentivizes greater utilization of zero-emission renewable energy, thus reducing curtailment. However, the TSCO framework’s ability to react to real-time conditions allows it to absorb significantly more renewable energy across all price points, maintaining a curtailment level that is less than half that of the DC-Frame. This again underscores the value of tight system integration for maximizing the use of clean energy resources.
VI Conclusion
This paper has addressed the critical and intertwined challenges of rising energy consumption in CPNs and the increasing instability of power grids due to high renewable energy penetration. We have argued that treating these two complex systems in a decoupled manner leads to suboptimal outcomes, characterized by high costs, significant carbon emissions, and wasted renewable energy. To overcome these limitations, a novel TSCO framework was proposed, designed to synergistically manage power system operations and CPN task scheduling. The results clearly demonstrate that by enabling the CPN to act as an active, flexible participant in grid operations, significant benefits can be realized. Compared to a decoupled approach, our integrated TSCO framework reduced total operational costs and carbon emissions. Most notably, it slashed renewable energy curtailment by over $60\%$ compared to conventional cost-only optimization, all while maintaining a high job success rate of over $98.5\%$ for computational tasks.
The findings of this research have significant practical implications for both power system operators and the owners of large-scale computational infrastructure. For grid operators, flexible CPNs can serve as a valuable source of ancillary services, reducing the need for expensive battery storage or fossil-fuel reserves. For CPN operators, participating in such programs can create new revenue streams and drastically reduce operational expenditures and carbon footprints.

### v2

@@
 Abstract
-The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which promise to deliver ubiquitous and on-demand computational resources.
-However, the immense energy consumption of these networks poses a significant sustainability challenge.
-Simultaneously, power grids are grappling with the instability introduced by the high penetration of intermittent renewable energy sources (RES).
+The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which deliver ubiquitous and on-demand computational resources.
+However, the immense energy consumption poses a significant sustainability challenge.
+Simultaneously, power grids are grappling with the instability by the high penetration of intermittent renewable energy sources (RES).
 This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations.
-The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment (SUC) stage and a real-time operational stage.
-The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a Deep Reinforcement Learning (DRL) agent.
-This agent makes intelligent, carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity.
+The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment stage and a real-time operational stage.
+The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a deep reinforcement learning agent.
+It makes carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity.
 Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches.
-Results demonstrate that the proposed framework reduces total carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining stringent Quality of Service (QoS) for computational tasks.
-Index Terms:
-Computing Power Network (CPN), renewable energy, data center, carbon-aware scheduling, deep reinforcement learning.
+It can reduce carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining a stringent quality of service for computational tasks.
 I Introduction
@@
 However, RES are inherently variable, non-dispatchable, and dependent on weather conditions.
-This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10], [11].
+This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10].
 Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict.
@@
 The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [11].
-This reframes the problem from ”how to power the CPN cleanly” to ”how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”.
+This reframes the problem from “how to power the CPN cleanly” to “how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”.
 This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
@@
   • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid.
-The model incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
+It incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
   • A Novel Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages.
@@
 II Related Works
-This section provides a structured review of the literature across four key domains that intersect in this work, Computing Power Network (CPN) architectures, integrated energy system management, carbon-aware workload scheduling, and optimization under uncertainty in power systems.
+This section provides a structured review of the literature across three key domains that intersect in this work, including Computing Power Network (CPN) architectures, integrated energy system management, and carbon-aware workload scheduling.
 This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
 II-A Computing Power Network Architectures and Scheduling
-The concept of the CPN has evolved from earlier paradigms like cloud, fog, and edge computing, with the primary goal of creating a unified network for ubiquitous computing resources.
+The CPN concept has evolved from earlier paradigms, such as cloud, fog, and edge computing, with the primary goal of creating a unified network for ubiquitous computing resources.
 Early research focused on defining the architecture and core features of CPNs, such as intent-driven operation, closed-loop autonomy, and elastic scheduling.
@@
 For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques.
-Other works have explored task offloading in terminal-side CPNs or the joint selection of routing paths and computing nodes [14].
+Other works have explored task offloading in terminal-side CPNs [13] or the joint selection of routing paths and computing nodes [14].
 While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored.
@@
 II-B Integrated Energy System Management
-In the power systems domain, the concept of coordinating generation, transmission, and consumption has been studied extensively under the ”Generation-Grid-Load-Storage” integrated operation model [11].
-This paradigm seeks to improve system safety, economy, and reliability through the coordinated interaction of all system components.
+In the power systems domain, the concept of coordinating generation, transmission, and consumption has been studied extensively under the “Generation-Grid-Load-Storage” integrated operation model [15].
+This paradigm seeks to improve system safety, economy, and reliability through the coordinated interaction of all components.
 Research in this area includes the development of multi-timescale optimal dispatching strategies [16], economic dispatch models that incorporate DR, and Optimal Power Flow (OPF) formulations that aim to minimize generation costs while respecting network constraints [17].
@@
 However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload.
-They treat the CPN as a ”black box” load, missing the opportunity to optimize its internal operations in concert with the grid.
+They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid [19].
 II-C Carbon-Aware Workload Scheduling
-The field of green computing has produced a significant body of work on carbon-aware scheduling for geo-distributed data centers.
-A common strategy is spatio-temporal scheduling [18], which involves shifting computational workloads in time (”temporal flexibility”) or space (”geographical flexibility”) to data centers with lower electricity prices or cleaner energy mixes.
+The green computing field has produced a substantial body of work on carbon-aware scheduling for geographically distributed data centers.
+A common strategy is spatio-temporal scheduling [20], which involves shifting computational workloads in time or space to data centers with lower electricity prices or cleaner energy mixes.
 These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
 To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [21].
-This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a ”virtual queue” that tracks the deviation from the budget.
+This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget.
 The scheduler is then penalized for actions that increase this queue length [22].
 While powerful, this line of research suffers from a critical limitation.
-It almost universally treats the power grid as an exogenous system.
+It almost universally treats the power grid as an exogenous system [23].
 The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions.
@@
 III System Architecture and Problem Formulation
-This section presents the mathematical models for the CPN and the integrated power system. These models form the basis of our co-optimization framework, as shown in Fig. 1.
+This section presents the mathematical models for the CPN and the integrated power system.
+These models form the basis of our co-optimization framework, as shown in Fig.
+1.
+For clarity, Table I summarizes the key notations used throughout this paper.
 Figure 1: CPN and power grid co-optimization architecture.
-III-A Computing Power Network (CPN) Model
+TABLE I: Summary of Key Notations
+Symbol Description
+$\mathcal{N}$ Set of CPN nodes (data centers)
+$\mathcal{G}_{C}$ Set of conventional thermal generators
+$\mathcal{G}_{R}$ Set of renewable energy (RES) generators (wind/solar)
+$\mathcal{J}$ Set of computational jobs in CPN
+$\mathcal{T}_{k}$ Set of subtasks of job $k$
+$\mathcal{I}$ Set of buses in the IEEE 30-bus power system
+$C_{n,h}^{comp}$ Computing capacity of hardware $h$ at CPN node $n$ (in FLOPS)
+$P_{n,t}^{CPN}$ Total power consumption of CPN node $n$ at time $t$
+$P_{g,t}$ Power output of conventional generator $g$ at time $t$ (in MWh)
+$P_{g,t,\omega}^{R,avail}$ Available power output of RES generator $g$ at time $t$ under scenario $\omega$
+$x_{k,\tau,n,h,t}$ Binary variable: 1 if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$
+$u_{g,t}$ Commitment status of conventional generator $g$ at time $t$ (1=on, 0=off)
+$SOC_{b,t,\omega}$ State-of-charge of battery energy storage system (BESS) $b$ at time $t$ under scenario $\omega$
+$\epsilon_{g}$ Carbon emission factor of generator $g$ (tons of $\text{CO}_{2}$ per MWh)
+$E_{budget}$ Long-term carbon emission budget for the planning horizon
+$LMP_{i,t}$ Locational marginal price at bus $i$ and time $t$ (in \$/MWh)
+$MCI_{i,t}$ Marginal carbon intensity at bus $i$ and time $t$
+$\Omega$ Set of uncertainty scenarios
+III-A Computing Power Network Model
 The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, representing data centers, and $\mathcal{L}$ is the set of communication links connecting them.
 III-A1 CPN Node Model
-Each CPN node $n\in\mathcal{N}$ is characterized by its available computational resources, such as CPUs, GPUs, and TPUs.
-For simplicity in this formulation, we aggregate these into a single measure of computing capacity, $C_{n}^{comp}$, expressed in Floating Point Operations Per Second (FLOPS).
-The power consumption of a CPN node is modeled as a linear function of its utilization [21], varying between an idle power $P_{n}^{idle}$ and a peak power $P_{n}^{peak}$ at full utilization.
-The total power consumed by node $n$ at time $t$ is a decision variable, $P_{n,t}^{CPN}$, determined by the scheduled tasks.
+Each CPN node $n\in N$ consists of heterogeneous hardware resources, denoted by the set $H=\{CPU,GPU,TPU\}$.
+For each hardware type $h\in H$, we define its key parameters: $C_{n,h}^{comp}$ is computing capacity in FLOPS); $P_{n,h}^{idle}$ is idle power consumption; $P_{n,h}^{peak}$ denotes peak power consumption at full utilization; $\alpha_{n,h},\beta_{n,h}$ are nonlinear coefficients calibrated from real-world hardware traces.
+The power consumption of hardware $h$ at node $n$ follows a quadratic nonlinear model, capturing the superlinear power growth of accelerators (e.g., GPU/TPU) under high utilization:
+$P_{n,h,t}^{comp}=P_{n,h}^{idle}+(\alpha_{n,h}\cdot u_{n,h,t}^{2}+\beta_{n,h}\cdot u_{n,h,t})\cdot(P_{n,h}^{peak}-P_{n,h}^{idle}),$ (1)
+where $u_{n,h,t}$ is the utilization of hardwar $h$ at node $n$ and time $t$, calculated as:
+$u_{n,h,t}=\frac{\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau,h}\cdot x_{k,\tau,n,h,t}}{C_{n,h}^{comp}},$ (2)
+where $r_{\tau,h}$ denotes the resource requirement of subtask $\tau$ for hardwarehh, set to 0 if $\tau$ does not require $h$.
+$x_{k,\tau,n,h,t}$ is a binary variable.
+Specifically, $1$ if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$, 0 otherwise.
+The total power consumption of nodenn and time $t$ is the sum of power from all hardware types:
+$P_{n,t}^{CPN}=\sum_{h\in H}P_{n,h,t}^{comp}.$ (3)
 III-A2 Task Model
@@
 Each job $k$ has an arrival time $A_{k}$ and a hard end-to-end deadline $D_{k}$.
+Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$, in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$.
+The execution time of subtask $\tau$ on hardware $h$ of node $n$ is:
+$t_{\tau,n,h}=\begin{cases}\frac{w_{\tau}}{C_{n,h}^{\mathrm{comp}}},&\text{if }h\in h_{\tau}\\ \infty,&\text{otherwise}\end{cases}\quad\forall n,h\in h_{\tau},\tau\in\mathcal{T}_{k}.$ (4)
+III-A3 Communication Link Model
+The communication network between CPN nodes is modeled as a directed graph $\mathcal{G}_{comm}=(\mathcal{N},\mathcal{L})$, where $\mathcal{L}$ denotes the set of bidirectional communication links between nodes.
+For each link from the source node to the destination node $(n_{src},n_{dest})\in\mathcal{L}$, two key parameters are defined based on real-world network traces:
+  • Bandwidth: $B_{n_{src},n_{dest}}$ (GB/s) is the maximum data transmission rate of the link, fixed for dedicated CPN networks.
+  • Latency parameters: It includes $lat_{n_{src},n_{dest}}^{base}$ (s, base latency without data transmission) and $\gamma_{n_{src},n_{dest}}$ (s/GB, dynamic latency coefficient related to data volume).
+Then, the total latency for transmitting a subtask $\tau$ with data volume $d_{\tau}$ from $n_{src}$ to $n_{dest}$ is:
+    $lat_{n_{src},n_{dest}}(\tau)=lat_{n_{src},n_{dest}}^{base}+\gamma_{n_{src},n_{dest}}\cdot d_{\tau}.$ (5)
+    The data migration time (i.e., the duration of data transmission) is determined by link bandwidth:
+    $t_{migrate,\tau,n_{src},n_{dest}}=\frac{d_{\tau}}{B_{n_{src},n_{dest}}}.$ (6)
 III-B Integrated Power System Model
@@
     Task Assignment: Each sub-task $\tau$ of each job $k$ must be scheduled exactly once.
-    $\sum_{n\in\mathcal{N}}\sum_{t=A_{k}}^{D_{k}}x_{k,\tau,n,t}=1\quad\forall k,\tau\in\mathcal{T}_{k}.$ (7)
+    $\sum_{n\in\mathcal{N}}\sum_{h\in\mathcal{h_{\tau}}}\sum_{t=A_{k}}^{D_{k}}x_{k,\tau,n,h,t}=1\quad\forall k,\tau\in\mathcal{T}_{k}.$ (13)
     Precedence Constraints: For any dependency $(\tau_{i},\tau_{j})\in\mathcal{E}_{k}$, the start time of $\tau_{j}$ must be after the finish time of $\tau_{i}$.
@@
     Node Resource Capacity: The total resource demand of tasks scheduled on node $n$ at time $t$ cannot exceed its capacity.
-    $\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau}x_{k,\tau,n,t}\leq C_{n}^{comp}\quad\forall n,t.$ (8)
-  • Coupling Constraint: This is the critical link. The power consumed by a CPN node $n$ (co-located at bus $i$) is determined by the tasks scheduled on it.
-    $P_{i,t,\omega}^{CPN}=P_{n}^{idle}+(P_{n}^{peak}-P_{n}^{idle})\frac{\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}}{C_{n}^{comp}}.$ (9)
-    This equation makes the CPN load term $P_{i,t,\omega}^{CPN}$ in the power balance equation an endogenous variable dependent on the scheduling decisions $x_{k,\tau,n,t}$.
+    $\sum_{k\in\mathcal{J}}\sum_{\tau\in\mathcal{T}_{k}}r_{\tau,h}x_{k,\tau,n,h,t}\leq C_{n,h}^{comp}\quad\forall n,h\in h_{\tau},t.$ (14)
+    Migration Transmission Time Constraint: For any subtask $\tau$ migrated from node $n_{src}$ to $n_{dest}$ (i.e., $x_{k,\tau,n_{src},h,t_{1}}=0$ and $x_{k,\tau,n_{dest},h,t_{2}}=1$ with $t_{2}>t_{1}$), the start time of $\tau$ at $n_{dest}$ must account for data migration time:
+    $start_{k,\tau,n_{dest},h,t_{2}}\geq$ $finish_{k,\tau_{pre},n_{src},h^{\prime},t_{1}}+$ (15)
+    $t_{migrate,\tau,n_{src},n_{dest}},$
+    $\forall(\tau_{pre},\tau)\in E_{k},h\in h_{\tau},h^{\prime}\in h_{\tau_{pre}},$ (16)
+    where $finish_{k,\tau_{pre},n_{src},h^{\prime},t_{1}}$ is the completion time of predecessor subtask $\tau_{pre}$ at $n_{src}$, and $t_{migrate,\tau,n_{src},n_{dest}}$ is calculated via Eq. (6).
+    Link Bandwidth Constraint: For each communication link $(n_{src},n_{dest})\in\mathcal{L}$, the total data transmission rate of all concurrent migration tasks must not exceed the link bandwidth:
+    $\sum_{\begin{subarray}{c}migrate(\tau,n_{src}\to n_{dest})\end{subarray}}\frac{d_{\tau}}{t_{migrate,\tau,n_{src},n_{dest}}}\leq B_{n_{src},n_{dest}},$ (17)
+    $\forall t,(n_{src},n_{dest})\in\mathcal{L},$ (18)
+    were $migrate(\tau,n_{src}\to n_{dest})$ denotes that subtask $\tau$ is migrated from $n_{src}$ to $n_{dest}$, and $\tau^{\prime}$ is the time interval of data transmission.
+  • Coupling Constraint: The total power consumption of CPN node $n$ co-located at bus $i$ is determined by the nonlinear power model of heterogeneous hardware and scheduling decisions:
+    $P_{i,n,t}^{CPN}=\sum_{h\in h_{\tau}}[P_{n,h}^{idle}+$ $(\alpha_{n,h}\cdot u_{n,h,t}^{2}+\beta_{n,h}\cdot u_{n,h,t})\cdot$ (19)
+    $(P_{n,h}^{peak}-P_{n,h}^{idle})].$
+    The CPN power consumption is a nonlinear function, while the real-time economic dispatch (ED) of the power grid is a linear programming problem.
+The direct coupling of these two will lead to a significant increase in the computational complexity of the optimization problem.
+To balance accuracy and tractability, we can use piecewise linearization to handle the nonlinear terms [30].
   • Carbon Budget Constraint: A long-term constraint on total carbon emissions is imposed to ensure sustainability goals are met.
@@
     1. 1.
-      Real-Time Economic Dispatch (ED): Solves for the optimal power output of the committed generators and the charge/discharge schedule of BESS to meet the actual load at minimum cost, while respecting all grid constraints.
+      Real-Time ED: Solves for the optimal power output of the committed generators and the charge/discharge schedule of BESS to meet the actual load at minimum cost, while respecting all grid constraints.
     2. 2.
@@
 Its decisions are informed by the real-time grid state, namely prices and carbon intensity, provided by the ED.
+The interface between the two stages is designed to ensure tight coupling while maintaining computational tractability.
+Specifically, Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope.
+In return, Stage 2 feeds back key real-time operational data to Stage 1 for iterative improvement.
+It includes the deviation between actual and forecasted RES generation, the CPN task completion rate and deadline adherence, and the cumulative carbon emissions up to the current time.
+This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem, e.g., adjusting the probability of high RES curtailment scenarios based on real-time absorption capacity.
+The interface operates on a daily rolling horizon.
+The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.
 This hierarchical structure allows the framework to be both economically optimal from a long-term planning perspective and highly adaptive to short-term dynamics.
@@
     1. 1.
-      CPN State: Characteristics of tasks in the queue (e.g., resource requirements, deadlines), current resource utilization and power consumption of each CPN node.
+      CPN State: Characteristics of tasks in the queue (e.g., resource requirements, deadlines), current resource and bandwidth utilization, and power consumption of each CPN node.
     2. 2.
@@
 These signals are provided by the real-time ED solution.
+    Furthermore, the state vector $s_{t}$ is augmented with two day-ahead constraint tracking metrics:
+    $Remaining_{n,h,t}=C_{n,h,t}^{comp,reserve}-\sum_{k,\tau}r_{\tau,h}x_{k,\tau,n,h,t-1},$ (23)
+    $Budget_{t}=E_{budget}-\sum_{\tau=1}^{t}\sum_{\omega\in\Omega}\pi_{\omega}E_{\tau,\omega},$ (24)
+    The former remains a computational resource reserve of node $n$ at time $t$.
+The latter remains the carbon budget for the planning horizon.
+These metrics enable the DRL agent to proactively avoid approaching constraint boundaries.
   • Action ($a_{t}\in\mathcal{A}$): For the task at the head of the queue, the agent selects an action from a discrete set.
-An action is a tuple $(n,\tau_{type})$ representing the decision to assign the task to node $n$ to be processed by resource type $\tau_{type}$ (e.g., CPU, GPU).
+An action is a tuple $(n,h,\tau_{type})$ representing the decision to assign the task to node $n$ to be processed by resource type $\tau_{type}$ (e.g., CPU, GPU).
 The action space also includes deferring the task.
+The action space is further constrained by the day-ahead resource reservations from Stage 1.
+For any task assignment action $(n,h,\tau_{type})$, the total resource demand of tasks scheduled on node $n$ at time $t$, i.e., $\sum_{k,\tau}r_{\tau,h}x_{k,\tau,n,h,t}$ must not exceed the day-ahead reserved capacity $C_{n,h,t}^{comp,reserve}$.
+This constraint is enforced directly in the DRL agent’s action selection process by masking infeasible assignments, e.g., nodes with remaining capacity $\leq$ task resource requirement $r_{\tau},h$.
   • Reward ($R_{t}\in\mathcal{R}$): The reward function is carefully designed to guide the agent towards the overall optimization objective. It is a weighted sum of multiple components:
@@
     $Q_{t+1}=\max(0,Q_{t}+\text{Carbon}_{t}-E_{budget}/T).$ (26)
-    The reward is then penalized by an additional term proportional to $Q_{t}\times\text{Carbon}_{t}$, which strongly discourages carbon-intensive actions when the system is already over its carbon budget.
+    The reward is then penalized by an additional term proportional to $Q_{t}\text{Carbon}_{t}$, which strongly discourages carbon-intensive actions when the system is already over its carbon budget.
 And the $\text{Penalty}_{t}$ represents a large negative penalty for missing a task’s deadline.
+The intuition behind this virtual carbon queue is to track the deviation between cumulative carbon emissions and the long-term budget.
+$Q_{t}$ increases when the current carbon emission $\text{Carbon}_{t}$ exceeds the average allowable emission $E_{budget}/T$ and resets to 0 if emissions are within the budget.
+When $Q_{t}$ grows, indicating the system is approaching or exceeding the carbon budget, the additional penalty term $Q_{t}\text{Carbon}_{t}$ in the reward function strongly discourages the DRL agent from assigning tasks to high-carbon-intensity nodes.
+It then ensures that short-term scheduling decisions do not violate the long-term sustainability goal while avoiding excessive sacrifice of economic efficiency.
+  • Constraint Adherence and Feedback Mechanism: To address potential deviations from the day-ahead plan, a two-tier feedback mechanism is implemented:
+    1. 1.
+      Real-Time Infeasibility Correction: If the real-time ED detects that the CPN power demand $P_{i,t}^{CPN}$ violates grid constraints, such as transmission line limits or generator ramp constraints, the ED module sends a “constraint violation signal” to the DRL agent.
+The agent then temporarily narrows its action space by increasing the penalty weight $w_{pen}$ for tasks assigned to nodes causing violations, or masking those nodes for 1-2 time steps until the grid state stabilizes.
+    2. 2.
+      Rolling Day-Ahead Plan Update: If the cumulative deviation between real-time CPN resource utilization and day-ahead reserve, i.e., $\sum_{t}|\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}-C_{n,t}^{comp,reserve}|$, exceeds a predefined threshold over 4 consecutive hours, the day-ahead SUC module is triggered to perform a mid-day rolling update.
+Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the latest real-time data.
 Given the large, continuous state space and discrete action space, a value-based DRL algorithm such as Deep Q-Network (DQN) or its advanced variants (e.g., Dueling DQN, Rainbow) is suitable [34].
@@
 Output : 24-hour unit commitment $\{u_{g,t}^{*}\}$, Real-time power dispatch, Real-time CPN task schedule
+1ex
 Stage 1: Day-Ahead SUC (solved once daily):
@@
 7. Obtain final 24-hour unit commitment schedule $\{u_{g,t}^{*}\}$
+1ex
 Stage 2: Real-Time Operation (for $t=1,\dots,T$):
@@
    5. Obtain real-time LMPs and MCIs for all CPN node locations
-   6. Construct state vector $s_{t}\leftarrow$ (CPN state + Grid state)
-   7. DRL agent takes action $a_{t}\leftarrow\pi(s_{t})$ to schedule next CPN task
+   6. Construct state vector $s_{t}\leftarrow$ (CPN state + Grid state + resource/carbon budget)
+   7. Mask infeasible actions, DRL agent takes action $a_{t}\leftarrow\pi(s_{t})$ to schedule task
    8. Update CPN power demand $P_{i,t}^{CPN}$ based on action $a_{t}$
-   9. Execute scheduled task and power dispatch
-   10. Update BESS SOC and CPN resource status
+   9. Correct grid constraint violations (if any), execute dispatch/scheduling
+   10. Update BESS SOC, CPN resource status and virtual carbon queue
 Algorithm 1 Two-Stage Co-Optimization (TSCO) Framework
@@
 The size of the master problem also grows with each iteration as Benders cuts are added.
-  • Subproblems: For each of the $|\Omega|$ scenarios, a linear program (LP) is solved.
+  • Subproblems: For each of the $\Omega$ scenarios, a linear program (LP) is solved.
 The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints.
@@
 A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS.
+Additional experiments on convergence, scalability, ablation, and computational burden are conducted to further verify the framework’s engineering feasibility and core value.
 V-A Simulation Setup
@@
 These datasets are used to generate 100 distinct 24-hour scenarios for the SUC problem.
-Figure 3: Baseline performance comparison. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
 The CPN consists of 5 geo-distributed nodes, each co-located with a major load bus in the IEEE 30-bus system.
@@
 Historical hourly locational marginal price (LMP) data is sourced from the California Independent System Operator (CAISO) public database [^9] https://www.gridstatus.io/live/caiso.
+To ensure the generality of the experimental results, all values are presented as the mean ± standard deviation derived from 10 independent runs.
+Figure 3: Baseline performance comparison. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
 V-B Comparison Schemes
@@
   • Cost-Only Optimizer (CO-Opt) [35]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero.
-This represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without explicit consideration for environmental impact.
+It represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without consideration for environmental impact.
   • Renewable-Greedy Scheduler (RG-Sched) [36]: This is a heuristic-based CPN scheduling approach where tasks are always dispatched to the CPN node with the highest instantaneous RES power availability.
@@
 This baseline tests the efficacy of a simple “follow the renewables” strategy that ignores grid constraints and economic signals.
-  • Decoupled Framework (DC-Frame) [34]: This baseline represents the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately.
+  • Decoupled Framework (DC-Frame) [37]: It is the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately.
 The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities.
@@
 This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
-V-C Simulation Results
-Figure 4: Sensitivity analysis with varying carbon price. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
+V-C Comparison Performances
 V-C1 Baseline Performance Comparison
-In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of \$50\\$50/ton.
-Each data point presented is the result of $10$ independent simulation runs, from which we calculate the mean and standard deviation and present in Fig.
+In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton.
+The mean and standard deviation are present in Fig.
 3.
@@
 In contrast, the RG-Sched baseline suffers from a poor job success rate $85.2\%$ and high tardiness because its singular focus on RES availability often leads it to schedule tasks on nodes that are already congested, highlighting the need for a holistic system view.
+Figure 4: Sensitivity analysis with varying carbon price. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
 V-C2 Sensitivity to Carbon Price
-To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from \$25\\$25/ton to \$150\\$150/ton.
+To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $\mathdollar 25$/ton to $\mathdollar 150$/ton.
 This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design.
@@
 This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points.
-The RG-Sched method is shown as fixed points, as their operational strategy does not adapt to the carbon price.
+The RG-Sched method is shown as a fixed point, as their operational strategy does not adapt to the carbon price.
+TABLE II: Convergence Performance Comparison
+Method Load Intensity Convergence Iterations Convergence Time (min) Bound Gap (\$)
+Direct-MILP Low - (infeasible) Larger than 360 (timeout) -
+Medium - (infeasible) Larger than 360 (timeout) -
+High - (infeasible) Larger than 360 (timeout) -
+Improved-Benders Low 28.3±2.1 45.7±3.2 128.5±15.3
+Medium 35.6±2.8 58.2±4.1 156.3±18.7
+High 42.1±3.5 72.4±5.6 189.7±21.2
+TSCO-Benders Low 21.5±1.8 32.4±2.5 98.6±12.4
+Medium 27.8±2.3 43.1±3.7 121.4±16.5
+High 33.2±2.9 55.8±4.8 145.2±19.3
 It also shows how RES curtailment is affected by the carbon price.
@@
 This again underscores the value of tight system integration for maximizing the use of clean energy resources.
+V-D Convergence Comparison
+To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders in [32].
+The results are summarized in Table II.
+Among them, the mid-range workload represents an average daily task arrival rate of $800$ jobs per day with a total computational workload of $5\times 10^{15}$ FLOPs, and an average resource requirement of $4$ CPU cores and 2 GPU cores or equivalent TPU resources per job.
+And the low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
+Direct-MILP fails to converge within 6 hours for all scenarios due to the high dimensionality of the SUC problem.
+Both Benders-based methods achieve convergence, but TSCO-Benders outperforms Improved-Benders by $24.2\%-27.9\%$ in convergence iterations and $29.1\%-31.3\%$ in convergence time.
+This is because the tailored Benders cuts in TSCO effectively reduce the search space of the master problem.
+The bound gap of TSCO-Benders is consistently below $\mathdollar 150$, indicating high solution accuracy.
+For medium load, TSCO-Benders takes an average of $43.1$ minutes to converge, which is acceptable for day-ahead planning.
+V-E Scalability Verification
+We also expand the system scale in three dimensions to test the scalability of the TSCO framework. The results are shown in Table III.
+TABLE III: Scalability Performance Under Different System Scales
+CPN Nodes RES Scenarios Gen. Count Day-Ahead Conv. Time (min) Day-Ahead Conv. Iterations Real-Time ED Time (s/step) DRL Inference Time (ms/task) Performance Retention Rate (%)
+5 100 6 43.1±3.7 27.8±2.3 1.2±0.1 8.5±0.6 100.0±0.0
+10 200 8 68.5±4.9 35.2±2.7 1.8±0.2 10.3±0.8 96.7±0.8
+15 300 10 95.3±6.2 42.7±3.1 2.5±0.3 12.1±1.0 93.5±1.1
+20 500 12 132.7±7.8 51.4±3.6 3.3±0.4 14.8±1.2 90.2±1.5
+Figure 5: Ablation test. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
+As the system scale expands with 20 nodes, 500 scenarios, and 12 generators, the day-ahead convergence time increases to 132.7 minutes.
+It is still within the 24-hour planning window, and the real-time ED/DRL time remains at the millisecond-to-second level.
+The performance retention rate is $90.2\%$, meaning the carbon emission reduction rate and RES utilization only decrease by $9.8\%$ compared to the baseline.
+This demonstrates that TSCO’s hierarchical structure effectively isolates computational complexity.
+The day-ahead stage tolerates longer computation for large-scale optimization, while the real-time stage maintains fast response via LP and DRL inference.
+The framework thus meets the requirements of large-scale CPN-grid integration.
+V-F Ablation Test
+To validate the necessity of key modules in the TSCO framework, we design four variants by removing core components one by one.
+In this test, we remove the nonlinear power consumption model of CPN from the TSCO and replace it with a simple linear model; Constraints of the communication link; Carbon-aware DRL; and integration with the two-stage process in TSCO.
+Its simulation parameters are set in accordance with those of Section V-C-1).
+The results are presented in Fig.
+5.
+Specifically, the nonlinear power model reduces carbon emissions by $7.1\%$ and operational cost by $3.0\%$ compared to the linear model, as it accurately captures GPU/TPU power dynamics and avoids over/underestimating CPN load.
+Meanwhile, communication constraints improve task success rate by $5.7\%$ and reduce average tardiness by $57.1\%$, as they prevent migration failures caused by bandwidth/latency bottlenecks.
+Moreover, the carbon-aware DRL reduces emissions by $28.3\%$ without significant economic loss, proving its ability to balance sustainability and cost.
+Furthermore, Two-stage synergy reduces operational cost by $8.4\%$ and RES curtailment rate by $55.6\%$, highlighting the value of real-time feedback to day-ahead planning.
+Therefore, we integrate the full framework of these modules to achieve optimal comprehensive performance.
+V-G Analysis of Computational Burden
+To verify the engineering feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB).
+The results are shown in Table IV.
+TABLE IV: Computational Burden Data
+Stage Metric Value (mins)
+Day-Ahead SUC Average Convergence Iterations 27.8±2.3
+Average Time per Iteration (min) 1.55±0.12
+Average Total Convergence Time (min) 43.1±3.7
+Master Problem Solving Time (min/iter) 0.32±0.05
+Subproblems Solving Time (min/iter, parallel) 1.23±0.10
+Real-Time Operation ED Average Solving Time (s/step) 1.2±0.1
+DRL Inference Time (ms/task) 8.5±0.6
+DRL Training Total Training Time (hours) 48.3±3.2
+The day-ahead SUC stage has an average convergence time of $43.1$ minutes, which is far less than the $24$-hour planning cycle, leaving sufficient time for scenario updates and human review.
+The real-time ED solves in $1.2$ seconds per step, namely a $5$-minute interval.
+Meanwhile, DRL infers in $8.5$ milliseconds per task, meeting the requirements of real-time scheduling with second-level response.
+The offline DRL training takes $48.3$ hours, which is acceptable as it only needs to be performed once before deployment with incremental fine-tuning every $3$ months based on new data.
+These quantitative results confirm that the TSCO framework balances optimization accuracy and computational efficiency, making it suitable for practical engineering deployment.
 VI Conclusion
@@
 Most notably, it slashed renewable energy curtailment by over $60\%$ compared to conventional cost-only optimization, all while maintaining a high job success rate of over $98.5\%$ for computational tasks.
-The findings of this research have significant practical implications for both power system operators and the owners of large-scale computational infrastructure.
-For grid operators, flexible CPNs can serve as a valuable source of ancillary services, reducing the need for expensive battery storage or fossil-fuel reserves.
-For CPN operators, participating in such programs can create new revenue streams and drastically reduce operational expenditures and carbon footprints.

### v3

@@
 Abstract
-The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which deliver ubiquitous and on-demand computational resources.
-However, the immense energy consumption poses a significant sustainability challenge.
-Simultaneously, power grids are grappling with the instability by the high penetration of intermittent renewable energy sources (RES).
-This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations.
+The proliferation of large-scale AI and data-intensive applications has driven the development of Computing Power Networks (CPN).
+It is a key paradigm for delivering ubiquitous, on-demand computational services with high efficiency.
+However, CPNs face dual challenges in service computing.
+Immense energy consumption threatens sustainable operations.
+And the integration with power grids also features high penetration of intermittent Renewable Energy Sources (RES), complicating task scheduling while ensuring Quality of Service (QoS).
+To address these issues, this paper proposes a novel Two-Stage Co-Optimization (TSCO) framework.
+It synergistically coordinates CPN task scheduling and power system dispatch, aiming to optimize service performance while achieving low-carbon operations.
 The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment stage and a real-time operational stage.
@@
 It makes carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity.
-Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches.
-It can reduce carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining a stringent quality of service for computational tasks.
+Extensive simulations demonstrate that the TSCO outperforms baseline approaches significantly.
+It reduces carbon emissions by $16.2\%$ and operational costs by $12.7\%$, while decreasing RES curtailment by over $60\%$, maintaining a task success rate of $98.5\%$, and minimizing average task tardiness to $12.3$s.
+This work advances cross-domain service optimization in CPNs.
 I Introduction
@@
 The digital transformation of the global economy is fueling an unprecedented demand for computational power.
-The rise of artificial intelligence (AI), big data analytics, and the Internet of Things (IoT) has pushed traditional centralized cloud computing architectures to their limits [1].
+The rise of artificial intelligence (AI) has pushed traditional centralized cloud computing architectures to their limits [1].
 In response, a new paradigm known as the Computing Power Network (CPN) has emerged [2].
-A CPN aims to interconnect vast, geographically distributed, and heterogeneous computing resources, spanning from large-scale data centers to edge nodes, into a unified, programmable fabric [3].
-The core vision of CPN is to break down resource silos and enable the flexible, on-demand scheduling and allocation of computing, storage, and network resources, thereby improving efficiency and enabling novel, low-latency applications.
-Concurrently, the global energy sector is undergoing a profound transition towards sustainability, marked by massive investments in Renewable Energy Sources (RES) such as wind and solar power [4].
-This shift presents a powerful, yet challenging, symbiotic relationship with the burgeoning CPN paradigm.
-On one hand, CPNs are voracious consumers of electricity; data centers alone are projected to account for a significant portion of global electricity load growth, with AI workloads being a primary driver.
+It aims to interconnect vast, distributed, and heterogeneous computing resources into a unified, programmable fabric [3].
+The core vision of CPN is to break down resource silos and enable the flexible, on-demand scheduling and allocation of computing, storage, and network resources, thereby enabling novel, low-latency applications.
+Unlike grid computing or conventional data centers, CPNs explicitly integrate computing task scheduling with power system dispatch.
+Concurrently, the global energy sector is undergoing a profound transition towards sustainability, marked by massive investments in Renewable Energy Sources (RES) and Battery Energy Storage Systems (BESS) [4].
+This shift presents a symbiotic relationship with the burgeoning CPN paradigm.
+On the one hand, CPNs are voracious consumers of electricity; data centers alone are projected to account for a significant portion of global electricity load growth, with AI workloads being a primary driver.
 Globally, data centers already account for $1-2\%$ of total electricity consumption, a figure comparable to the aviation industry.
 Driven by the explosive growth of AI, this demand is projected to double by 2030 [5].
-In the United States, data center electricity usage is forecast to climb from $4.4\%$ of the national total in 2023 to as high as $12\%$ by 2028111https://www.energy.gov/articles/doe-releases-new-report-evaluating-increase-electricity-demand-data-centers, with AI workloads shifting from a minor component to a primary driver of this expansion.
+In the United States, data center electricity usage is forecast to climb from $4.4\%$ of the national total in 2023 to as high as $12\%$ by 2028111https://www.energy.gov/articles/doe-releases-new-report-evaluating-increase-electricity-demand-data-centers, with AI workloads becoming a primary driver of this expansion.
 The availability of clean, renewable energy offers a direct path to decarbonize this massive computational infrastructure[6].
@@
 I-B Research Motivation
-The central challenge lies in a fundamental misalignment between the operational dynamics of CPNs and RES-dominated power grids.
-CPNs are designed to host a wide array of computational tasks, from latency-sensitive services like cloud virtual reality to compute-intensive, batch-processing jobs like large model training [8], [9].
+The core challenge lies in a misalignment between the operational dynamics of CPNs and RES-dominated power grids.
+CPNs are designed to host a wide array of computational tasks, from latency-sensitive services, to compute-intensive, batch-processing jobs [8].
 These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS).
-However, RES are inherently variable, non-dispatchable, and dependent on weather conditions.
-This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10].
+However, RES are inherently variable and dependent on weather conditions.
+This intermittency introduces significant volatility into the power grid.
+It often results in the wasteful curtailment of clean energy when generation exceeds demand [9].
 Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict.
@@
 The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset.
-The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [11].
+The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [10], [11].
 This reframes the problem from “how to power the CPN cleanly” to “how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”.
@@
 I-C Contributions of the Paper
-This paper proposes a holistic, hierarchical co-optimization framework that treats the power system and the CPN as a single, integrated entity.
+In the present work, we propose a holistic, hierarchical co-optimization framework that jointly optimizes power system operations and CPN task scheduling.
 By breaking down the silos between power system operators and CPN schedulers, the framework unlocks significant economic and environmental benefits.
 It coordinates power generation dispatch with computational task scheduling across multiple timescales to align CPN energy demand with the availability of low-cost, low-carbon renewable energy.
-As far as we know, this is one of the first efforts to jointly optimize the power grid and the CPN.
-The main contributions of this work are summarized as follows:
+To the best of our knowledge, this is one of the first efforts to jointly optimize the power grid and the CPN.
+The main contributions of this work are summarized next:
   • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid.
-It incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
-  • A Novel Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages.
+It incorporates conventional thermal generators, stochastic RES, and BESS, while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
+  • A Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages.
 A day-ahead planning stage addresses long-term unit commitment and resource reservation, while a real-time operational stage handles dynamic economic dispatch and adaptive task scheduling.
@@
 This technique effectively decouples the integer commitment decisions from the continuous dispatch variables, ensuring the problem remains computationally tractable even for large systems and numerous uncertainty scenarios.
-  • Adaptive Real-Time Scheduling with Deep Reinforcement Learning (DRL): A DRL agent is developed for the real-time CPN task scheduling subproblem.
-This model-free approach enables fast, adaptive, and carbon-aware scheduling decisions in response to the highly dynamic and complex state of the joint CPN-grid system, complementing the model-based optimization of the power dispatch.
-  • Extensive High-Fidelity Performance Evaluation: The efficacy of the proposed TSCO framework is validated through extensive simulations using realistic data.
-The evaluation demonstrates significant improvements in carbon emissions, operational costs, and RES utilization when compared against decoupled and carbon-agnostic baseline strategies.
+  • Adaptive Real-Time Scheduling with Deep Reinforcement Learning (DRL): A DRL agent is developed for the real-time CPN task scheduling.
+This model-free approach enables fast, adaptive, and carbon-aware scheduling decisions in response to the highly dynamic and complex state of the joint CPN-grid system.
 I-D Paper Structure
-The remainder of this paper is structured as follows.
-Section II provides a critical review of related work in CPN scheduling, integrated energy systems, and carbon-aware computing.
+The remainder of the paper is structured as follows.
+Section II provides a review of related work in CPN scheduling, integrated energy systems, and carbon-aware computing.
 Section III presents the detailed mathematical formulation of the CPN and power system models, defining the joint optimization problem.
@@
 II Related Works
-This section provides a structured review of the literature across three key domains that intersect in this work, including Computing Power Network (CPN) architectures, integrated energy system management, and carbon-aware workload scheduling.
-This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
 II-A Computing Power Network Architectures and Scheduling
@@
 The primary objectives have traditionally been to optimize QoS metrics.
-For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques.
-Other works have explored task offloading in terminal-side CPNs [13] or the joint selection of routing paths and computing nodes [14].
+For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques [13].
+Other works have explored task offloading in terminal-side CPNs or the joint selection of routing paths and computing nodes [14].
+Two other works also considered the collaboration between the CPN and the power grid.
+Wen et al.
+[15], [16] proposed a spatiotemporal task scheduling scheme for CPNs, leveraging a customized DRL algorithm to optimize energy consumption and carbon emissions within CPNs.
+However, their work focuses solely on intra-domain optimization, treating energy supply as an exogenous input without involving power system dynamics.
+Zhong et al.
+[17] integrated Electric Vehicles (EVs) into CPN and Power Distribution Network (PDN) coordination.
+They use EVs as dual carriers of energy storage and computing resources for task offloading.
+Nevertheless, their CPN scheduling is subordinate to cost minimization of EV-PDN synergy, lacking deep coupling with bulk power systems and carbon-aware optimization.
+In contrast, our TSCO framework emphasizes bidirectional coupling between CPN and bulk power systems, treating CPN as a flexible load to enhance grid stability and low-carbon operations.
+Table I compares the approaches with ours.
 While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored.
-Existing CPN scheduling literature typically treats the power grid as an external, static entity, overlooking the potential for dynamic, symbiotic interaction.
+TABLE I: Comparison of CPN and Power Grid Coordination
+Key Dimension TSCO (Our Work) [15], [16] [17]
+Research Paradigm Cross-domain co-optimization Intra-CPN fine-grained scheduling Multi-agent synergy
+Collaborative Entities CPN, bulk power system CPN CPN, PDN, EVs
+Optimization Core Low carbon, RES curtailment reduction, grid stability, communication QoS Energy efficiency, task delay, load balancing System cost minimization
+System Interaction Bidirectional (CPN and grid) No interaction Unidirectional (CPN/EV constrained by PDN)
+Key Innovation Bidirectional coupling; endogenous carbon intensity Triple-selection (nodes/paths/time); intra-CPN coordination EV dual-resource (energy/computing) exploitation
 II-B Integrated Energy System Management
@@
 However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload.
-They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid [19].
+They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid.
 II-C Carbon-Aware Workload Scheduling
-The green computing field has produced a substantial body of work on carbon-aware scheduling for geographically distributed data centers.
+A substantial body of work on carbon-aware scheduling for geographically distributed data centers.
 A common strategy is spatio-temporal scheduling [22], which involves shifting computational workloads in time or space to data centers with lower electricity prices or cleaner energy mixes.
-These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
+These methods often leverage real-time carbon intensity to guide scheduling.
 To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [23].
-This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget.
+This technique transforms a long-term average constraint into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget.
 The scheduler is then penalized for actions that increase this queue length [24].
-While powerful, this line of research suffers from a critical limitation.
-It almost universally treats the power grid as an exogenous system [23].
+This line of research suffers from a critical limitation.
+It almost universally treats the power grid as an exogenous system.
 The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions.
-This assumption breaks down at scale, as the collective actions of large CPNs can and will influence grid operations, market prices, and the generation mix, thereby altering the very signals they are responding to.
+This assumption breaks down at scale, as the collective actions of large CPNs can influence grid operations, market prices, and the generation mix.
+Specifically, we have done a study comparing four countries: Germany, the Netherlands, France, and Brazil, as these have very different energy mixes and therefore the variability of the carbon intensity is very different among them.
 II-D Research Gap Summary
@@
   • Simplifies the power grid: CPN and carbon-aware scheduling studies treat the grid as a static source of price and carbon signals, ignoring the feedback loop where scheduling decisions impact the grid.
-  • Simplifies the CPN: Power system and DR studies model flexible loads like data centers in an overly simplistic manner, failing to capture the rich internal complexity of CPN workloads, resource heterogeneity, and QoS constraints.
-This leads to a new class of control problem where the decision variables are distributed across two separate domains.
+  • Simplifies the CPN: Power system and DR studies model flexible loads in an overly simplistic manner, failing to capture the rich internal complexity of CPN workloads, resource heterogeneity, and QoS constraints.
+This leads to a new class of control problems where the decision variables are distributed across two separate domains.
 The system dynamics are characterized by a mix of well-understood physics (the power grid) and complex, stochastic behavior (the CPN).
-A model-based optimization approach is ill-suited for the fast, dynamic CPN scheduling, while a model-free AI approach cannot guarantee adherence to the hard physical constraints of the power grid.
-Consequently, there is a clear need for a hybrid framework that endogenously models the bi-directional interactions between the two systems.
-Our TSCO framework, which combines large-scale, model-based optimization for slow, physics-heavy planning with a model-free, adaptive AI technique for fast, complex real-time scheduling, is designed specifically to fill this gap.
+A model-based approach is ill-suited for fast, dynamic CPN scheduling, while a model-free AI approach cannot guarantee adherence to hard physical constraints of the power grid.
+Consequently, there is a need for a novel framework that endogenously models the bidirectional interactions between them.
+Our TSCO framework, which combines large-scale, model-based optimization for slow, physics-heavy planning with a model-free, adaptive AI technique for fast, complex real-time scheduling, is designed specifically.
 III System Architecture and Problem Formulation
-This section presents the mathematical models for the CPN and the integrated power system.
-These models form the basis of our co-optimization framework, as shown in Fig.
-1.
-For clarity, Table I summarizes the key notations used throughout this paper.
+These models form the basis of our co-optimization framework, as shown in Fig. 1.
 Figure 1: CPN and power grid co-optimization architecture.
-TABLE I: Summary of Key Notations
-Symbol Description
-$\mathcal{N}$ Set of CPN nodes (data centers)
-$\mathcal{G}_{C}$ Set of conventional thermal generators
-$\mathcal{G}_{R}$ Set of renewable energy (RES) generators (wind/solar)
-$\mathcal{J}$ Set of computational jobs in CPN
-$\mathcal{T}_{k}$ Set of subtasks of job $k$
-$\mathcal{I}$ Set of buses in the IEEE 30-bus power system
-$C_{n,h}^{comp}$ Computing capacity of hardware $h$ at CPN node $n$ (in FLOPS)
-$P_{n,t}^{CPN}$ Total power consumption of CPN node $n$ at time $t$
-$P_{g,t}$ Power output of conventional generator $g$ at time $t$ (in MWh)
-$P_{g,t,\omega}^{R,avail}$ Available power output of RES generator $g$ at time $t$ under scenario $\omega$
-$x_{k,\tau,n,h,t}$ Binary variable: 1 if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$
-$u_{g,t}$ Commitment status of conventional generator $g$ at time $t$ (1=on, 0=off)
-$SOC_{b,t,\omega}$ State-of-charge of battery energy storage system (BESS) $b$ at time $t$ under scenario $\omega$
-$\epsilon_{g}$ Carbon emission factor of generator $g$ (tons of $\text{CO}_{2}$ per MWh)
-$E_{budget}$ Long-term carbon emission budget for the planning horizon
-$LMP_{i,t}$ Locational marginal price at bus $i$ and time $t$ (in \$/MWh)
-$MCI_{i,t}$ Marginal carbon intensity at bus $i$ and time $t$
-$\Omega$ Set of uncertainty scenarios
 III-A Computing Power Network Model
-The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, representing data centers, and $\mathcal{L}$ is the set of communication links connecting them.
+The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, and $\mathcal{L}$ is the set of communication links connecting them.
 III-A1 CPN Node Model
@@
 Each job $k$ has an arrival time $A_{k}$ and a hard end-to-end deadline $D_{k}$.
-Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$, in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$.
+Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$ in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$.
 The execution time of subtask $\tau$ on hardware $h$ of node $n$ is:
@@
 III-B Integrated Power System Model
-The power grid is modeled based on the IEEE 30-bus test system [^3] https://icseg.iti.illinois.edu/ieee-30-bus-system/, a standard benchmark for power system studies.
-The system consists of a set of buses $\mathcal{I}$ connected by transmission lines.
-Each CPN node $n\in\mathcal{N}$ is co-located with a specific load bus $i\in\mathcal{I}$.
+The system consists of a set of buses $\mathcal{I}$ connected by transmission lines. Each CPN node $n\in\mathcal{N}$ is co-located with a specific load bus $i\in\mathcal{I}$.
 III-B1 Conventional Generation
@@
 The environmental impact is quantified through carbon emissions.
-The carbon intensity of the grid is not static but depends on the real-time generation mix.
+The carbon intensity is not static but depends on the real-time generation mix.
 The total carbon emission rate at time $t$ in scenario $\omega$, $E_{t,\omega}$, is calculated as the sum of emissions from all active generators:
@@
 This endogenous calculation is crucial, as it directly links dispatch decisions to carbon output.
-We also leverage real-world marginal carbon intensity data, such as that provided by WattTime, to inform the real-time DRL agent about the emissions impact of consuming an additional unit of electricity at a specific location and time.
+We also leverage real-world marginal carbon intensity data, such as that provided by WattTime [^2] https://watttime.org/data-science/data-signals/, which informs the real-time DRL agent about the emission impact resulting from an additional unit of electricity consumption at a specific location and time.
 III-D Joint Optimization Problem Formulation
 The overarching goal is to co-optimize the power system operation and CPN task scheduling to minimize the total expected system cost over a planning horizon $T$.
-The total cost comprises the operational costs of the power system and the monetized cost of carbon emissions.
-The problem is formulated as a large-scale, two-stage stochastic mixed-integer linear program (MILP).
+It comprises the operational costs of the power system and the monetized cost of carbon emissions.
+It is formulated as a large-scale, two-stage stochastic mixed-integer linear program (MILP).
 Objective Function:
@@
 The objective minimizes the expected sum of three components across all scenarios: (1) the quadratic fuel costs of conventional generators, (2) the costs associated with starting up ($SU_{g}$) and shutting down ($SD_{g}$) these generators based on their commitment status $u_{g,t}$, and (3) a carbon tax, where $\lambda_{CO2}$ is the price of carbon and $E_{t,\omega}$ is the total emissions.
-Key Constraints: The optimization is subject to a comprehensive set of constraints that couple the two systems:
+The optimization is subject to a comprehensive set of constraints that couple the two systems:
   • Power System Constraints (for each $t,\omega$):
@@
     $\sum_{g\in\mathcal{G}(i)}P_{g,t,\omega}+\sum_{b\in\mathcal{B}(i)}(P_{b,t,\omega}^{dis}-P_{b,t,\omega}^{chg})\\ -(P_{i,t}^{D}+P_{i,t,\omega}^{CPN})=\sum_{j\in\mathcal{I}}B_{ij}(\theta_{i,t,\omega}-\theta_{j,t,\omega}),$ (11)
-    where, $\mathcal{G}(i)$ and $\mathcal{B}(i)$ are generators and BESS at bus $i$, $P_{i,t,\omega}^{CPN}$ is the CPN power demand at that bus, and the right-hand side represents the net power flow out of the bus.
-    Transmission Line Limits: The power flow $F_{ij}$ on each line $(i,j)$ must not exceed its thermal limit $F_{ij}^{max}$.
+    where $\mathcal{G}(i)$ and $\mathcal{B}(i)$ are generators and BESS at bus $i$. $P_{i,t,\omega}^{CPN}$ is the CPN power demand. The right side represents the net power flow out of the bus.
+    Transmission Line Constraints: The power flow $F_{ij}$ on each line $(i,j)$ must not exceed its thermal limit $F_{ij}^{max}$.
     $F_{ij}^{max}\leq B_{ij}(\theta_{i,t,\omega}-\theta_{j,t,\omega})\leq F_{ij}^{max}.$ (12)
@@
 IV Two-Stage Co-Optimization (TSCO) Framework
-The joint optimization problem formulated in Section III is a large-scale, non-convex, mixed-integer stochastic program, which is computationally intractable to solve directly [31].
-To address this challenge, we propose a Two-Stage Co-Optimization (TSCO) framework that decomposes the problem by decision timescale and complexity.
-The framework combines model-based optimization for long-term, system-wide planning with a model-free, AI-based approach for fast, adaptive real-time control, as shown in Fig.
+The joint optimization problem formulated in Section III is a large-scale, non-convex, mixed-integer stochastic program, whose direct solution is computationally prohibitive at realistic scales [32].
+Our TSCO framework combines model-based optimization for long-term, system-wide planning with a model-free, AI-based approach for fast, adaptive real-time control, as shown in Fig.
 2.
@@
 IV-A Hierarchical Stochastic Optimization Structure
-The TSCO framework decomposes the problem into two distinct stages, reflecting the natural hierarchy of power system operations:
+The TSCO framework decomposes the problem into two distinct parts, reflecting the power system’s natural hierarchy.
   • Stage 1 (Day-Ahead Planning): This stage solves an SUC problem for the upcoming 24-hour horizon.
-It makes the ”here-and-now” decisions, which are binding across all potential future scenarios.
+It makes the “here-and-now” decisions, which are binding across all potential future scenarios.
 These decisions include the commitment (on/off) status $u_{g,t}$ of conventional generators and high-level energy reservation for BESS and CPN workload classes.
@@
 This stage is computationally intensive but is performed only once per day.
-  • Stage 2 (Real-Time Operation): This stage operates at a much faster timescale (e.g., 5-15 minute intervals) and makes ”wait-and-see” recourse decisions as uncertainty unfolds.
+  • Stage 2 (Real-Time Operation): This stage operates at a much faster timescale (e.g., 5-15 minute intervals) and makes “wait-and-see” recourse decisions as uncertainty unfolds.
 As the actual RES generation and CPN task arrivals are revealed, this stage executes two parallel, tightly coupled processes:
@@
 The interface between the two stages is designed to ensure tight coupling while maintaining computational tractability.
-Specifically, Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope.
+Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope.
 In return, Stage 2 feeds back key real-time operational data to Stage 1 for iterative improvement.
 It includes the deviation between actual and forecasted RES generation, the CPN task completion rate and deadline adherence, and the cumulative carbon emissions up to the current time.
-This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem, e.g., adjusting the probability of high RES curtailment scenarios based on real-time absorption capacity.
-The interface operates on a daily rolling horizon.
-The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.
-This hierarchical structure allows the framework to be both economically optimal from a long-term planning perspective and highly adaptive to short-term dynamics.
+This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem.The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.
 IV-B Benders Decomposition for the Day-Ahead SUC
@@
       Rolling Day-Ahead Plan Update: If the cumulative deviation between real-time CPN resource utilization and day-ahead reserve, i.e., $\sum_{t}|\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}-C_{n,t}^{comp,reserve}|$, exceeds a predefined threshold over 4 consecutive hours, the day-ahead SUC module is triggered to perform a mid-day rolling update.
-Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the latest real-time data.
+Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the real-time data.
 Given the large, continuous state space and discrete action space, a value-based DRL algorithm such as Deep Q-Network (DQN) or its advanced variants (e.g., Dueling DQN, Rainbow) is suitable [35].
@@
 Algorithm 1 Two-Stage Co-Optimization (TSCO) Framework
-Additionally, the computational complexity of the TSCO framework is best analyzed by examining its two stages separately.
+Additionally, the computational complexity of the TSCO framework is best analyzed by examining its two stages.
 IV-D1 Stage 1: Day-Ahead SUC
@@
 The size of the master problem also grows with each iteration as Benders cuts are added.
-  • Subproblems: For each of the $\Omega$ scenarios, a linear program (LP) is solved.
+  • Subproblems: For each $\Omega$ scenario, a linear program (LP) is solved.
 The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints.
@@
 V Performance Evaluation
-This section presents a comprehensive empirical validation of the proposed Two-Stage Co-Optimization (TSCO) framework.
 A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS.
@@
 For solving the MILP and LP problems in the Benders decomposition, we use the Gurobi optimizer.
-The simulation is based on a modified IEEE 30-bus test system.
+The simulation is based on a modified IEEE 30-bus test system [^3] https://icseg.iti.illinois.edu/ieee-30-bus-system/.
 It includes 6 conventional thermal generators, 4 utility-scale BESS units, and 5 large-scale renewable generation sites (3 solar, 2 wind).To ensure realism, we use real-world time-series data to model the stochastic RES generation.
@@
   • Cost-Only Optimizer (CO-Opt) [36]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero.
-It represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without consideration for environmental impact.
+It represents the current industry-standard approach of economic dispatch.
   • Renewable-Greedy Scheduler (RG-Sched) [37]: This is a heuristic-based CPN scheduling approach where tasks are always dispatched to the CPN node with the highest instantaneous RES power availability.
-The power system is not co-optimized; it simply reacts to the resulting CPN load profile.
-This baseline tests the efficacy of a simple “follow the renewables” strategy that ignores grid constraints and economic signals.
   • Decoupled Framework (DC-Frame) [17]: It is the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately.
@@
 Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals.
-This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
+  • DRL-based Spatiotemporal Scheduler (ST-Sched) [15], [16]: This baseline adopts a customized DRL approach for intra-CPN fine-grained scheduling.
+It features a triple-selection action space to optimize energy consumption, carbon emissions, and task delay while ensuring load balancing.
 V-C Comparison Performances
 V-C1 Baseline Performance Comparison
-In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton.
+In this scenario, all five methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton.
 The mean and standard deviation are present in Fig.
@@
 Its reliance on static, day-ahead signals prevents it from adapting to real-time deviations between forecasted and actual grid conditions, leading to $12.7\%$ higher costs and $16.2\%$ higher emissions.
+The ST-Sched focuses on intra-CPN fine-grained scheduling via customized DRL.
+It demonstrates competitive performance relative to DC-Frame and CO-Opt but falls short of TSCO.
+It achieves $8.4\%$ higher operational costs and $4.4\%$ higher carbon emissions than TSCO.
+Its lack of cross-domain coordination with the power system prevents it from leveraging real-time grid dynamics (e.g., surplus renewable energy) to optimize overall efficiency.
 A key finding is TSCO’s ability to significantly reduce RES curtailment.
-By treating the CPN as a flexible load that can absorb surplus renewable generation in real-time, TSCO reduces curtailment by over $60\%$ compared to DC-Frame and CO-Opt.
+By treating the CPN as a flexible load that can absorb surplus renewable generation in real-time, TSCO reduces curtailment by over $60\%$ compared to DC-Frame and CO-Opt, and by $37.1\%$ compared to ST-Sched.
 This demonstrates the value of co-optimization in turning a major energy consumer into a valuable grid-stabilizing asset.
@@
 V-C2 Sensitivity to Carbon Price
-To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $\mathdollar 25$/ton to $\mathdollar 150$/ton.
-This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design.
+To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $25$ to 150150/ton.
+This analysis focuses on the TSCO, DC-Frame, and ST-Sched methods, as the other two baselines are insensitive to carbon price by design.
 The comprehensive results are presented in Fig.
 4.
-The TSCO framework consistently achieves a better trade-off.
-For any given carbon emission level, it operates at a lower cost than the DC-Frame.
-As the carbon price increases, both methods are incentivized to reduce emissions, but TSCO does so more efficiently.
-This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points.
-The RG-Sched method is shown as a fixed point, as their operational strategy does not adapt to the carbon price.
+TSCO achieves a superior economic-environmental trade-off across all metrics at every carbon price point.
+All three carbon-aware methods cut emissions with rising carbon prices, but TSCO does so with the highest efficiency at $19.3\%$.
+Economically, TSCO also demonstrates optimal performance.
+Its operational cost rises by the smallest margin with increasing carbon prices, remaining approximately $7\%$ lower than ST-Sched and $14\%$ lower than DC-Frame across all price points.
+ST-Sched’s single-domain intra-CPN DRL optimization delivers moderate cost and emission improvements over DC-Frame but falls short of TSCO.
+It lacks cross-domain coordination to leverage grid-side low-carbon, low-cost resources.
+DC-Frame’s decoupled design, reliant on static day-ahead grid signals, fails to adapt to real-time deviations.
+It leads to the highest operational costs and the least efficient emission reduction among them.
+For RES curtailment, rising carbon prices drive greater renewable utilization for all methods, with TSCO recording the largest reduction at $14.3\%$.
+Its curtailment levels are $37.1\%$ lower than ST-Sched and less than half that of DC-Frame.
+This underscores TSCO’s unique ability to treat the CPN as a flexible load to absorb surplus grid-side renewable energy.
 TABLE II: Convergence Performance Comparison
@@
 High 33.2±2.9 55.8±4.8 145.2±19.3
-It also shows how RES curtailment is affected by the carbon price.
-For both carbon-aware methods, increasing the price of carbon incentivizes greater utilization of zero-emission renewable energy, thus reducing curtailment.
-However, the TSCO framework’s ability to react to real-time conditions allows it to absorb significantly more renewable energy across all price points, maintaining a curtailment level that is less than half that of the DC-Frame.
-This again underscores the value of tight system integration for maximizing the use of clean energy resources.
+Notably, TSCO’s environmental and economic gains do not compromise CPN service quality.
+It retains the highest job success rate and lowest average tardiness across all carbon prices.
+ST-Sched also delivers strong QoS performance due to its refined intra-CPN load balancing.
+While DC-Frame experiences noticeable QoS degradation with rising carbon prices due to its rigid precomputed grid signals failing to account for real-time CPN congestion.
 V-D Convergence Comparison
-To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders in [32].
+To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders [33].
 The results are summarized in Table II.
 Among them, the mid-range workload represents an average daily task arrival rate of $800$ jobs per day with a total computational workload of $5\times 10^{15}$ FLOPs, and an average resource requirement of $4$ CPU cores and $2$ GPU cores or equivalent TPU resources per job.
-And the low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
+The low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
 Direct-MILP fails to converge within 6 hours for all scenarios due to the high dimensionality of the SUC problem.
@@
 V-G Analysis of Computational Burden
-To verify the engineering feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB).
+To verify the feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB).
 The results are shown in Table IV.
@@
 VI Conclusion
-This paper has addressed the critical and intertwined challenges of rising energy consumption in CPNs and the increasing instability of power grids due to high renewable energy penetration.
-We have argued that treating these two complex systems in a decoupled manner leads to suboptimal outcomes, characterized by high costs, significant carbon emissions, and wasted renewable energy.
-To overcome these limitations, a novel TSCO framework was proposed, designed to synergistically manage power system operations and CPN task scheduling.
-The results clearly demonstrate that by enabling the CPN to act as an active, flexible participant in grid operations, significant benefits can be realized.
-Compared to a decoupled approach, our integrated TSCO framework reduced total operational costs and carbon emissions.
-Most notably, it slashed renewable energy curtailment by over $60\%$ compared to conventional cost-only optimization, all while maintaining a high job success rate of over $98.5\%$ for computational tasks.
+This paper addresses the intertwined challenges of growing energy consumption in CPNs and power grid instability caused by high RES penetration.
+These two issues are critical for service computing sustainability and QoS.
+To overcome these limitations, we propose a novel TSCO framework that synergistically coordinates CPN task scheduling and power system dispatch.
+By decomposing the complex problem into a day-ahead stochastic unit commitment stage and a real-time operational stage, the framework aligns service QoS requirements with grid sustainability goals.
+Extensive simulations validate the TSCO framework’s superiority over baselines.
+It achieves a $16.2\%$ cut in carbon emissions and a $12.7\%$ reduction in operational costs, while cutting RES curtailment by more than $60\%$.
+Notably, these environmental and economic gains are achieved without compromising service reliability.
+The framework sustains a $98.5\%$ task success rate and limits average task tardiness to $12.3$s.
+This work advances cross-domain service optimization in CPNs, laying the groundwork for efficient, low-carbon service computing.